# Can LLMs Handle WebShell Detection? Overcoming Detection Challenges with Behavioral Function-Aware Framework

### v4

@@ Abstract
WebShell attacks, where malicious scripts are injected into web servers, pose a significant cybersecurity threat. Traditional machine learning and deep learning methods are often hampered by challenges such as the need for extensive training data, catastrophic forgetting, and poor generalization. Recently, Large Language Models (LLMs) have emerged as a powerful alternative for code-related tasks, but their potential in WebShell detection remains underexplored. In this paper, we make two major contributions: (1) a comprehensive evaluation of seven LLMs, including GPT-4, LLaMA 3.1 70B, and Qwen 2.5 variants, benchmarked against traditional sequence- and graph-based methods using a dataset of 26.59K PHP scripts, and (2) the Behavioral Function-Aware Detection (BFAD) framework, designed to address the specific challenges of applying LLMs to this domain. Our framework integrates three components: a Critical Function Filter that isolates malicious PHP function calls, a Context-Aware Code Extraction strategy that captures the most behaviorally indicative code segments, and Weighted Behavioral Function Profiling (WBFP) that enhances in-context learning by prioritizing the most relevant demonstrations based on discriminative function-level profiles. Our results show that, stemming from their distinct analytical strategies, larger LLMs achieve near-perfect precision but lower recall, while smaller models exhibit the opposite trade-off. However, all baseline models lag behind previous State-Of-The-Art (SOTA) methods. With the application of BFAD, the performance of all LLMs improves significantly, yielding an average F1 score increase of 13.82%. Notably, larger models like GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B now outperform SOTA benchmarks, while smaller models such as Qwen-2.5-Coder-3B achieve performance competitive with traditional methods. This work is the first to explore the feasibility and limitations of LLMs for WebShell detection and provides solutions to address the challenges in this task.
@@ 1 Introduction
The rapid growth of web applications and the expansion of cloud-based services have intensified the cybersecurity threat landscape, with WebShells emerging as a significant concern. WebShells are malicious scripts injected into web servers that allow attackers to remotely execute arbitrary commands, steal sensitive data, and compromise system integrity (Starov et al., 2016; Cui et al., 2018). According to a Cisco Talos report (Talos, 2024), threat groups used a variety of web shells against vulnerable or unpatched web applications in 35% of incidents in Q4 2024, a sharp increase from the previous quarter, when such activity was observed in only 10% of incidents. These attacks are particularly dangerous due to the stealthy nature of WebShells, which are rapidly evolving to evade traditional detection methods (Hannousse & Yahiouche, 2021).
In response to this challenge, the community has proposed several approaches. Rule-based methods that rely on predefined signatures or heuristics are increasingly ineffective against the complexity and diversity of modern WebShells (Le et al., 2021; Jinping et al., 2020). Machine learning models, especially deep learning techniques (Pu et al., 2022), have shown promise in addressing these threats. However, they require extensive training on large datasets–resources that are often difficult to obtain and sensitive in nature (Shang et al., 2024). In addition, these models face challenges such as catastrophic forgetting and poor generalization, especially when dealing with obfuscated or encrypted attacks (Jinping et al., 2020; Zhang et al., 2025).
Recently, Large Language Models (LLMs) have gained attention for a variety of code-related tasks, including code generation (Ma et al., 2024) and vulnerability detection (Liu & He, 2023; Wang et al., 2025). Studies have shown that with effective prompt engineering, LLMs can perform remarkably well without additional training (Nong et al., 2024; Trad & Chehab, 2025). They also provide interpretable explanations for their decisions, which is critical in cybersecurity contexts (Ma et al., 2024). Despite these advantages, there has been limited exploration of LLMs for WebShell detection.
Detecting WebShells with LLMs presents unique challenges that differ from other code analysis tasks. WebShells often employ obfuscation and encryption techniques and are embedded in large codebases dominated by benign content (Liu & He, 2023). For example, the largest WebShell in our dataset spans 1,386,438 tokens, far exceeding the context window of most LLMs, which risks truncating critical malicious segments when processing entire source files (Wang et al., 2025; Ceka et al., 2024). In addition, in-context learning (ICL) struggles in this domain: the variability and obfuscation of WebShells complicate the selection of effective demonstrations, and these examples further occupy significant context space, reducing capacity for the target code (Yuan et al., 2024). While recent research has focused on increasing the context length of LLMs (Chen et al., 2023), studies suggest that the performance of an LLM tends to degrade with longer inputs, and the low processing speed may also become unacceptable for practical use (Ma et al., 2024; Fang et al., 2024).
In this paper, we present two key contributions to advance the application of LLMs in WebShell detection.
First, we systematically evaluate LLMs in the context of WebShell detection, comparing their performance with traditional state-of-the-art (SOTA) machine learning methods. Specifically, we evaluate seven closed-source and open-source LLMs of different sizes–including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024)–on a dataset containing 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples). Our analysis reveals several key findings:
  • Larger LLMs, such as GPT-4 and Qwen 2.5 Coder 14B, achieve near-perfect precision but struggle with recall (e.g., GPT-4’s recall is 85.98%). Their deep contextual analysis, while powerful, can be deceived by well-disguised malicious code, leading to false negatives.
  • Smaller LLMs, such as Qwen 2.5 Coder 3B and Qwen 2.5 0.5B, exhibit high recall but suffer from low precision (e.g., Qwen 2.5 Coder 3B’s precision is 38.93%). This stems from their reliance on surface-level pattern matching, which often flags any use of sensitive functions without sufficient contextual validation.
  • Randomly selected ICL demonstrations degrade LLM detection performance; examples selected based on semantic similarity to the source code do not yield significant improvements.
For this comparison, we benchmark against several traditional methods, including GloVe+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023). The best-performing LLM, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, although it still lags behind GAT-based methods, which achieve an F1 score of 98.87%. [^1] While traditional models such as GAT have demonstrated superior performance, they require extensive training and significant data collection efforts. In contrast, LLMs enable direct detection through prompts, leveraging their pre-trained knowledge with minimal task-specific resources to deliver competitive results.
Second, we present the Behavioral Function-Aware Detection (BFAD) framework to address the identified challenges and improve the performance of LLMs to meet the requirements of downstream applications by achieving a more balanced trade-off between precision and recall. Our framework combines risk-filtering techniques with an enhanced ICL strategy that uses weighted demonstration selection to prioritize examples most closely related to key malicious behaviors. Experimental results show that our approach improves the average F1 score across all LLMs by 13.82%, with GPT-4 and Qwen 2.5 0.5B improving by 6.89% and 51.23%, respectively. In several cases, such as GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B, our approach enables performance that is competitive with or even superior to traditional methods.
To the best of our knowledge, this is the first work to systematically analyze the feasibility and limitations of applying LLMs to WebShell detection.
@@
2 Behavioral Function-Aware Detection Framework
We present the Behavioral Function-Aware Detection (BFAD) framework, a comprehensive solution designed to improve WebShell detection by identifying critical code segments and improving the quality of ICL demonstration selection. The architecture of BFAD, as shown in Figure 1, consists of three primary components: (a) Critical Function Filter, which is based on malicious behavior patterns and filters critical PHP function calls to identify key functions; (b) Context-Aware Code Extraction, which addresses the limitations of LLMs in handling long input sequences by selectively extracting critical code regions; and (c) Weighted Behavioral Function Profiling, which improves demonstration selection by calculating a weighted similarity score based on function-level profiling.
@@ 2.1 Critical Function Filter
WebShells typically rely on specific PHP functions that facilitate malicious actions such as code execution, data exfiltration, or obfuscation. However, these functions are often embedded in complex, obfuscated code, making detection difficult. To address this, we are developing a Critical Function Filter that classifies PHP functions into six different behavioral categories: Program Execution, Code Execution, Callback Functions, Network Communication, Information Gathering, and Obfuscation and Encryption. This taxonomy reflects the different roles these functions play in WebShells.
Specifically, the Program Execution category includes functions such as exec and system, which execute system-level commands and are often exploited in WebShells to execute arbitrary payloads. Similarly, Code Execution functions such as eval and preg_replace interpret input as executable code, allowing attackers to inject arbitrary scripts. The Callback Functions category includes functions such as array_map and register_shutdown_function, which allow dynamic invocation of functions often used to obfuscate malicious code.
In addition, Network Communication functions such as fsockopen and curl_init allow remote communication for data exfiltration and command-and-control operations. Information Gathering functions, such as phpinfo and getenv, are used by attackers to gather system details. Finally, Obfuscation and Encryption functions, such as base64_encode and openssl_encrypt, help disguise or encrypt malicious payloads to avoid detection.
Our statistical analysis (detailed in Appendix C) shows that WebShell files use critical functions far more often than benign files. On average, WebShells contain 22.76 calls to critical functions, compared to only 0.74 in benign files, underscoring their behavioral complexity and distinctiveness. [^2] While this disparity in critical function usage is significant, it alone does not reliably distinguish WebShells from benign files, since legitimate scripts may also call these functions. Therefore, we use LLMs for deeper contextual analysis to improve detection accuracy.
@@ 2.2 Context-Aware Code Extraction
Building on the Critical Function Filter, we introduce a Context-Aware Code Extraction strategy that identifies and extracts the critical code regions that indicate malicious behavior. These regions focus on the identified critical functions and their surrounding contexts, ensuring that the LLM focuses on the most relevant parts of the code.
@@
We reduce the input size by selectively extracting critical regions of code and merging overlapping segments while preserving behavioral specificity. However, this approach may inadvertently exclude certain global contextual information and increase the emphasis on critical functions, potentially leading to false positives when analyzing benign files that legitimately use such functions. To mitigate this, we append truncated, non-overlapping code segments when context length allows, ensuring that the model receives a balanced representation of local and global code context.
@@ 2.3 Weighted Behavioral Function Profiling
Based on the extracted critical code regions, we propose Weighted Behavioral Function Profiling (WBFP), a method that computes a weighted similarity score to identify behaviorally similar examples for ICL effectively. WBFP assigns weights to each function type based on its prevalence and usage in WebShell versus benign files, quantified by three metrics: coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$). The coverage difference measures the proportion of files containing a specific function across the two datasets. The frequency ratio is the ratio of the average number of occurrences of the function per file in WebShell files to the average number per file in benign files. The usage ratio reflects the total number of function occurrences in WebShell files compared to those in benign files. These metrics are combined to calculate a discrimination score for each function type
@@
where $\alpha$, $\beta$, and $\gamma$ are empirically determined weights, set to 1 for balanced contribution in our experiments (see Section 3.2). We normalize the discrimination scores to weights:
@@
WBFP then uses the embeddings $E(\cdot)$ generated by st-codesearch-distilroberta-base (Abi Akl, 2023; Al-Kaswan et al., 2023) to compute the similarity between the files $x$ and $y$. Let $\mathcal{F}$ denote all critical function types. For each function type $f\in\mathcal{F}$, we concatenate critical regions $R_{f}(x)$ and compute their embeddings:
$\mathbf{e}_{f}(x)=E(\text{concat}_{f}(x)),$
The semantic similarity between the files $x$ and $y$ for the function type $f$ is given by
@@
This weighted similarity prioritizes function types critical to WebShell behavior, reducing the impact of irrelevant semantic features. Using this score, WBFP ensures that ICL demonstrations accurately capture malicious patterns, improving detection performance.

### v5

@@ Abstract
WebShell attacks—where adversaries implant malicious scripts on web servers—remain a persistent threat. Prior machine-learning and deep-learning detectors typically depend on task-specific supervision and can be brittle under data scarcity, rapid concept drift, and out-of-distribution (OOD) deployment. Large language models (LLMs) have recently shown strong code understanding capabilities, but their reliability for WebShell detection remains unclear. We address this gap by (i) systematically evaluating seven LLMs (including GPT-4, LLaMA-3.1-70B, and Qwen-2.5 variants) against representative sequence- and graph-based baselines on 26.59K PHP scripts, and (ii) proposing Behavioral Function-Aware Detection (BFAD), a behavior-centric framework that adapts LLM inference to WebShell-specific execution patterns. BFAD anchors analysis on security-sensitive PHP functions via a Critical Function Filter, constructs compact LLM inputs with Context-Aware Code Extraction, and selects in-context demonstrations using Weighted Behavioral Function Profiling (WBFP), which ranks examples by a behavior-weighted, function-level similarity. Empirically, we observe a consistent precision–recall asymmetry: larger LLMs often achieve high precision but miss attacks (lower recall), while smaller models exhibit the opposite tendency; moreover, off-the-shelf LLM prompting underperforms established detectors. BFAD substantially improves all evaluated LLMs, boosting F1 by 13.82% on average; notably, GPT-4, LLaMA-3.1-70B, and Qwen-2.5-Coder-14B exceed prior SOTA benchmarks, while Qwen-2.5-Coder-3B becomes competitive with traditional methods. Overall, our results clarify when LLMs succeed or fail on WebShell detection, provide a practical recipe and highlight future directions for making LLM-based detection more reliable.
@@ 1 Introduction
The rapid growth of web applications and cloud services has expanded the attack surface of modern systems, making WebShells a persistent threat. A WebShell is a malicious script implanted on a web server that enables remote command execution, data exfiltration, and broader system compromise (Starov et al., 2016; Cui et al., 2018; Han, 2026). A Cisco Talos report (Talos, 2024) highlights the prevalence of such activity: web shells were observed in 35% of incidents in Q4 2024, up from 10% in the previous quarter. WebShells are particularly difficult to defend against because they are continually adapted—often via obfuscation or encryption—to evade conventional detection pipelines (Hannousse and Yahiouche, 2021).
To counter this threat, prior work spans signature/heuristic defenses and learned detectors. Rule-based approaches are increasingly brittle given the diversity and rapid evolution of WebShell variants (Le et al., 2021; Jinping et al., 2020). Learning-based methods, including deep models (Pu et al., 2022), can capture richer patterns but typically require substantial labeled data and costly retraining—resources that are often scarce or sensitive in practice (Shang et al., 2024). Moreover, these models can suffer from catastrophic forgetting under continual updates and may generalize poorly to novel obfuscation or encryption strategies (Jinping et al., 2020; Zhang et al., 2025).
Large language models (LLMs) have recently demonstrated strong performance on code-centric tasks, from generation to program understanding (Ma et al., 2024), and have been explored for security applications such as vulnerability detection (Liu and He, 2023; Wang et al., 2025; Han et al., 2026b). With appropriate prompting, LLMs can be adapted to new tasks without additional training (Nong et al., 2024; Trad and Chehab, 2025; Han et al., 2026a), and they can provide natural-language rationales that support analyst workflows (Ma et al., 2024). However, their effectiveness for WebShell detection remains insufficiently characterized.
Applying LLMs to WebShell detection introduces challenges that are less pronounced in standard code analysis. WebShell payloads are often heavily obfuscated or encrypted and may be buried within predominantly benign code (Liu and He, 2023). Naively feeding an entire file to an LLM can therefore be ineffective: in our dataset, the longest WebShell contains 1,386,438 tokens, far beyond typical context windows, so truncation can easily remove the malicious core (Wang et al., 2025; Ceka et al., 2024). In addition, in-context learning (ICL) is fragile in this domain: the diversity of obfuscation patterns complicates demonstration selection, and demonstrations themselves consume context budget that would otherwise be allocated to the target code (Yuan et al., 2024). While longer-context LLMs are actively studied (Chen et al., 2023), prior work suggests that performance can degrade on longer inputs and latency may become prohibitive in practice (Ma et al., 2024; Fang et al., 2024).
In this paper, we take a step toward making LLM-based WebShell detection both measurable and reliable.
First, we conduct a systematic evaluation of LLMs for WebShell detection and compare them with representative learned detectors. Concretely, we evaluate seven closed- and open-source LLMs spanning a wide range of scales—including GPT-4 (Achiam et al., 2023), LLaMA 3.1 70B (Grattafiori et al., 2024), Qwen 2.5 Coder (14B/3B) (Yang et al., 2024), and Qwen 2.5 (3B/1.5B/0.5B) (Yang et al., 2024)—on a dataset of 26.59K PHP scripts (4.93K WebShells and 21.66K benign samples). For this comparison, we benchmark against several traditional methods, including GloVe+SVM (Petridis, 2024; Rigutini et al., 2024), CodeBERT+Random Forest (Alghamdi et al., 2022), and graph-based approaches such as GAT (Kang et al., 2023).
Our experiments yield three main observations:
  • Scale shifts the error mode: larger LLMs (e.g., GPT-4, Qwen 2.5 Coder 14B) achieve very high precision but lower recall (e.g., GPT-4 recall 85.98%), whereas smaller models often trade precision for recall (e.g., Qwen 2.5 Coder 3B precision 38.93%).
  • Naive ICL is unreliable: randomly selected demonstrations can degrade performance, and selecting demonstrations by semantic similarity alone yields limited improvements.
  • LLMs do not close the gap by default: without task-specific adaptation, off-the-shelf LLM prompting remains behind strong learned detectors in overall F1. For example, the best-performing LLM baseline, Qwen 2.5 Coder 14B, achieves an F1 score of 96.39%, but still trails the GAT-based detector (98.87% F1). [^1] While traditional models such as GAT can achieve higher accuracy, they typically require task-specific training and sustained data collection. In contrast, LLMs can be deployed via prompting, leveraging pretraining with minimal task-specific supervision.
Second, we propose Behavioral Function-Aware Detection (BFAD), a behavior-centric framework designed to make LLM inference effective under these constraints and to better balance precision and recall. BFAD couples risk-aware filtering with an enhanced ICL strategy that weights demonstrations by their alignment with discriminative malicious behaviors. Across all evaluated LLMs, BFAD improves F1 by 13.82% on average; GPT-4 and Qwen 2.5 0.5B improve by 6.89% and 51.23%, respectively. For several models (including GPT-4, LLaMA 3.1 70B, Qwen 2.5 Coder 14B, and Qwen 2.5 Coder 3B), BFAD yields performance competitive with—and in some cases surpassing—strong traditional baselines.
Finally, we outline several forward-looking research directions suggested by our analysis, including (i) constructing privacy-preserving, large-scale synthetic benchmarks for stress-testing generalization, (ii) integrating graph-based behavioral representations with LLMs via multimodal alignment (e.g., graph encoders plus lightweight adapters), and (iii) designing agentic “fast–slow” detection pipelines with autonomous update loops to sustain performance under distribution shift.
To the best of our knowledge, this is the first work to systematically characterize when and why LLMs succeed or fail on WebShell detection, and to provide a practical framework that closes much of the gap to learned state-of-the-art detectors.
2 Related Work
WebShell Detection Techniques. Early WebShell detection systems were predominantly rule-based, relying on signatures and handcrafted heuristics to flag malicious scripts (Le et al., 2021; Jinping et al., 2020). While effective against known families, such approaches are brittle under obfuscation and rapid variant evolution because they depend on predefined patterns (Hannousse and Yahiouche, 2021). Subsequent machine-learning (ML) methods introduced feature-based classification using textual or behavioral signals. For instance, Jinping et al. (2020) combined Random Forest and CNN models with N-gram and TF-IDF features, reporting strong performance on PHP WebShells but noting sensitivity to dataset balance and encrypted samples. Deep learning has further expanded the design space by leveraging semantic and structural representations, including transformer-based embeddings (e.g., CodeBERT) (Pu et al., 2022) and graph neural networks for code structure modeling. Zhang et al. (2025) proposed MMFDetect, which fuses CodeBERT-CL semantics with CNN-extracted visual features from RGB-mapped PHP code. Despite steady progress, supervised ML/DL detectors typically require substantial labeled data (often scarce in security settings) and can generalize poorly to heavily obfuscated or previously unseen WebShell variants, while incurring non-trivial training and maintenance costs (Shang et al., 2024; Jinping et al., 2020).
LLMs for Code and Security Analysis. Large language models (LLMs) trained on massive code corpora have become a powerful interface for code understanding and generation, enabling applications ranging from synthesis (Ma et al., 2024) to vulnerability detection (Sun et al., 2024) and reliability assessment (Liu et al., 2024). In security contexts, LLMs are attractive because they support zero-shot and few-shot generalization via prompting and can provide natural-language rationales that improve analyst interpretability (Nong et al., 2024; Han et al., 2026c). For example, Ma et al. (2024) showed that LLMs can generate evasive WebShells with carefully designed prompts, and Sun et al. (2024) proposed retrieval-enhanced prompting for vulnerability reasoning. However, compared with these broader code-security applications, LLM-based WebShell detection remains relatively underexplored, especially for long, obfuscated scripts where malicious behavior may be sparsely embedded.
Challenges in Applying LLMs to WebShell Detection. Directly applying LLMs to WebShell detection faces two practical bottlenecks. First, long WebShell files can exceed fixed context windows, causing truncation that may omit the truly malicious region (Ceka et al., 2024; Wang et al., 2025). Moreover, performance can degrade as input length grows; Fang et al. (2024) reported substantial accuracy drops for obfuscated long-context code. Techniques such as chunking and long-context model architectures (e.g., sparse attention) can partially alleviate this issue (Guo et al., 2023; Wang et al., 2024b), but may sacrifice global context and introduce fragmentation (Wang et al., 2025). Second, in-context learning (ICL) is itself context-expensive: demonstrations consume a significant fraction of the prompt budget (Min et al., 2022; Wang et al., 2025), and naive retrieval (random or whole-file semantic similarity) often fails to capture the behavioral structure of WebShells. While information-theoretic criteria have been explored for ICL selection (Liu and Wang, 2023), existing formulations primarily target natural language classification and do not directly address behavior-centric code security tasks.
Taken together, the limitations of (i) rule-based rigidity, (ii) supervised ML/DL dependence on labeled data and weak OOD robustness, and (iii) LLM context and ICL constraints motivate BFAD. BFAD addresses long-context inputs via hybrid, behavior-anchored extraction and improves ICL reliability with WBFP, which retrieves demonstrations based on behavior-weighted, function-level profiles rather than generic similarity.
3 Behavioral Function-Aware Detection Framework
We present Behavioral Function-Aware Detection (BFAD), a framework that improves WebShell detection by (i) isolating behavior-relevant code snippets for LLM analysis and (ii) selecting in-context learning (ICL) demonstrations that better match the target file’s malicious behaviors. As illustrated in Figure 1, BFAD comprises three components: (a) a Critical Function Filter that maps PHP function calls to a behavior taxonomy and extracts behavior-relevant functions; (b) Context-Aware Code Extraction that constructs a compact yet informative code view under LLM context limits; and (c) Weighted Behavioral Function Profiling that ranks candidate demonstrations using a behavior-weighted similarity score computed from function-centric representations.
@@ 3.1 Critical Function Filter
WebShells often implement malicious behaviors—e.g., command execution, data exfiltration, and payload obfuscation—through a small set of high-risk PHP functions. In practice, these calls are frequently buried in long, noisy, and obfuscated code, which dilutes LLM attention and complicates robust detection. BFAD therefore begins with a Critical Function Filter that organizes PHP functions into six behavior categories: Program Execution, Code Execution, Callback Functions, Network Communication, Information Gathering, and Obfuscation and Encryption. This taxonomy captures the distinct roles that functions play in typical WebShell workflows.
Concretely, Program Execution includes functions such as exec and system that invoke system commands; Code Execution includes eval and preg_replace (with the /e modifier in legacy PHP) that can interpret strings as code; and Callback Functions includes array_map and register_shutdown_function that enable indirect invocation and are frequently used to hide control flow.
Network Communication includes fsockopen and curl_init for remote communication (e.g., command-and-control or exfiltration); Information Gathering includes phpinfo and getenv for environment reconnaissance; and Obfuscation and Encryption includes base64_encode and openssl_encrypt for disguising or encrypting payloads.
Our statistical analysis (Appendix B) indicates that WebShell files invoke these critical functions far more frequently than benign files: on average, WebShells contain 22.76 critical-function calls, compared to 0.74 in benign files. This gap suggests strong behavioral signals, but it is insufficient by itself because benign programs may legitimately use the same APIs. We therefore treat critical functions as anchors to locate behavior-relevant code and rely on LLMs to disambiguate benign versus malicious intent from context.
@@ 3.2 Context-Aware Code Extraction
Building on the Critical Function Filter, BFAD performs Context-Aware Code Extraction to construct an LLM-friendly representation of each file. The key idea is to center extraction around critical function occurrences and include their surrounding context, so that the LLM receives the minimal code necessary to infer the intent of each call.
@@
We reduce input length by extracting windows around critical calls and merging overlapping windows, which preserves local behavioral evidence while avoiding redundant tokens. This operation can remove global context and over-emphasize critical functions, potentially increasing false positives for benign scripts that legitimately call high-risk APIs. To mitigate this, when the context budget allows we append additional truncated, non-overlapping segments from the remaining code, providing complementary global signals without exceeding the model’s context limit.
@@ 3.3 Weighted Behavioral Function Profiling
Given extracted regions, BFAD selects ICL demonstrations via Weighted Behavioral Function Profiling (WBFP). WBFP ranks candidate examples by a behavior-weighted similarity that emphasizes function types that best discriminate WebShells from benign files. Specifically, for each critical function type $f$, we compute three statistics from the WebShell and benign corpora: coverage difference ($r_{c}$), frequency ratio ($r_{f}$), and usage ratio ($r_{u}$). Coverage difference measures how often $f$ appears at least once in a file across the two corpora; frequency ratio measures the ratio of the average per-file call count; and usage ratio measures the ratio of the total call counts. We combine them into a discrimination score:
@@
where $\alpha$, $\beta$, and $\gamma$ are scalar hyperparameters (set to 1 in our experiments for equal contribution; see Section 4.2). We normalize scores into weights:
@@
WBFP then computes similarity using embeddings $E(\cdot)$ from st-codesearch-distilroberta-base (Abi Akl, 2023; Al-Kaswan et al., 2023). Let $\mathcal{F}$ denote the set of critical function types. For each $f\in\mathcal{F}$, we concatenate the extracted regions associated with $f$ (denoted $R_{f}(x)$) and embed them as
$\mathbf{e}_{f}(x)=E(\text{concat}_{f}(x)).$
The function-type similarity between files $x$ and $y$ is computed by cosine similarity:
@@
This formulation prioritizes function types that are more indicative of WebShell behavior and down-weights less informative semantics, yielding demonstrations that better match the target file’s behavioral profile.